OpenStudy (anonymous):

Evaluate (2.25)^1/2??

OpenStudy (saifoo.khan):

\[\Huge \sqrt{2.25}= \frac32\]

OpenStudy (anonymous):

and???/

OpenStudy (saifoo.khan):

that's the answer. or write, 1.5

OpenStudy (anonymous):

1.5???????

OpenStudy (saifoo.khan):

3/2 = 1.5

OpenStudy (anonymous):

oh you converted the 2.25 into a fraction

OpenStudy (saifoo.khan):

not really.

OpenStudy (anonymous):

that is what i just did, I converted 2.25 into a mixed number and then converted that into a fraction, what did you do?

OpenStudy (saifoo.khan):

well, i used my calc. but the manual way is: 2.25 = 225/100 = 9/4 (9/4)^1/2 \[\sqrt{\frac94}\]=3/2

OpenStudy (anonymous):

what way is best?

OpenStudy (saifoo.khan):

mine.

OpenStudy (saifoo.khan):

the best of the best is using a Calculator.